satellite73 (satellite73):

what is wrong with this question? The equation \(y^2+4y+4x^2+4x+21=0\) can be changed into a vertex equation by completing the square. Which conic section can be made from the graph of this equation? A. ellipse B. hyperbola C. circle D. parabola

OpenStudy (anonymous):

Calculate the Discriminant. It is an ellipse.

OpenStudy (anonymous):

no actually it isn't

OpenStudy (anonymous):

Ok...

OpenStudy (anonymous):

question was just asked here it was i guess supposed to be an ellipse, but if you complete the square you see it is nothing

OpenStudy (anonymous):

You are trying to recover it in the standard form of ellipse \[\frac{ x^2 }{a^2 } + \frac{ y^2 }{ b^2 } = 1\] ?

OpenStudy (anonymous):

If you calculate its determinant, it comes out to be positive. This is the condition for a degenerate ellipse. So there are no real points to plot it. I don't know how wolframalpha engine is plotting it ! I think it is treating x as a variable and y as a parameter.

OpenStudy (anonymous):

How is that plotting a 3D structure ?

OpenStudy (anonymous):

Oh! In wolframalpha engine you need to search for y^2+4y+4x^2+4x+21 = 0, which is an equation. You have put an expression y^2+4y+4x^2+4x+21 there.